OpenStudy (anonymous):

If a and b are positive integers such that ab = 125, then (a-b)^(a+b-4) is equal to ?

OpenStudy (anonymous):

ab = 125 125 = 5*25 or 1*125 (no other integers multiply into 125) Therefore: \[f(a,b) = (a-b)^{a+b-4} \] \[f(125,1) = 2.49715552 × 10^{255}\] \[f(1,125) = 2.49715552 × 10^{255}\] \[f(25, 5) = 1.49011612 × 10^{18}\] \[f(5,25) = 1.49011612 × 10^{18}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

but

OpenStudy (anonymous):

options are: 1) 16 2) 25 3) 28 4) 30

OpenStudy (anonymous):

Did you write the question correctly?